OpenStudy (anonymous):

A bag contains 4 white and 6 black marbles. 28 students each reach in, select a marble, record its color, then replace the marble. a. What is the probability that exactly 10 students select white marbles? b. What is the probability that no more than 10 students select white marbles?

OpenStudy (zarkon):

binomial distribution

OpenStudy (anonymous):

?

OpenStudy (kropot72):

\[P(10/white)=\frac{28!}{10!18!}(0.4)^{10}(0.6)^{18}\]

OpenStudy (zarkon):

they want "no more than 10" so they will need to do a little more work.

OpenStudy (anonymous):

true, been stuck on this for ages...

OpenStudy (kropot72):

Let P(10 white) be the probability that 10 students pick white. Probability that no more than 10 pick white is [1 - P(10 white)]. @uj1024 What do you get for P(10 white) ?

OpenStudy (zarkon):

that is not quite right

OpenStudy (anonymous):

9/14?

OpenStudy (kropot72):

I get 0.14 for the probability that 10 will pick white using the equation that I posted earlier.

OpenStudy (anonymous):

UPDATE!!! i added question a. as a hint, although it has already been solved

OpenStudy (zarkon):

\[\sum_{x=0}^{10}{28\choose x}(0.4)^{x}(0.6)^{28-x}\]

OpenStudy (anonymous):

i dont understand the difference between exactly and no more than

OpenStudy (anonymous):

i dont think sequences and sigma notation have anything to do with this...

OpenStudy (zarkon):

no more than 10...means 10 is the max...so you can have anywhere between 0 and 10 inclusive

OpenStudy (anonymous):

so, 10! ?

OpenStudy (zarkon):

I suggest you read up on the binomial distribution

OpenStudy (anonymous):

is that related to binomial theorem? i havent i learnt about binomial distribution

OpenStudy (zarkon):

yes

OpenStudy (kropot72):

@uj1024 What Zarkon has said is right. Please read up on the binomial distribution and then you should be able to understand the solution that he has given.

OpenStudy (anonymous):

For each student, let the probability of picking a white marble be p, and the probablity of picking a black marble be q. (Note that q = 1 - p by definition.) The probability that the sequence picked will be wwbwwbwwbwwbwwbwwbwwbwwbwwb is \[p^{19} q^9\] Because white was drawn 19 times and black 9 times. Of course, this also the probality of drawing bwwbwwbwwbwwbwwbwwbwwbwwbww, or indeed any sequence that involves 19 whites and 9 blacks. How many such sequences are there? The number of ways of arranging 28 objects, 19 of which are identical (white) and another 9 of which are identical (black), or 28!/(19! 9!). The probability of getting any one of these sequences is the sum of all terms, or: \[\frac{28!}{19! 9!} p^{19} q^9\] We can immediately generalize to say the probability of getting any of the possible sequences of k whites and n-k blacks is: \[\frac{ n! }{k! (n-k)!} p^k q^{n-k}\] What you need is the probability that at least 18 students pick black marbles, so you need the sum of this expression from k = 0 (no white marbles) to k = n - 18 (all but 18 marbles white): \[\sum_{k=0}^{n-18} \frac{n!}{k! (n-k)!} p^k q^{n-k}\] Unfortunately, I know of no closed expression for this object. If I write a program to calculate this sum, for p = 4/10, n = 28, using the GNU GMP program, I get: \[\frac{1486137203415528633}{1490116119384765625}\approx 0.997330\]

OpenStudy (anonymous):

Oops, I typed in the wrong arguments (28 and 18) to my program. For n = 28 and kmax = 10 I get: \[\frac{2969550422392190877}{7450580596923828125}\approx 0.398566\] which seems just a bit more reasonable. Also, this is the GNU GMP library.

OpenStudy (zarkon):

as I wrote above \[\sum_{x=0}^{10}{28\choose x}(0.4)^{x}(0.6)^{28-x}\] which equals \[\frac{2969550422392190877}{7450580596923828125}\] :)

OpenStudy (anonymous):

Yes, I agree. I gave the longer explanation so that the OP could see WHY this is the correct solution. How did you get the numerical value, by the way?

OpenStudy (zarkon):

I just used my calculator

OpenStudy (anonymous):

Goodness, they make remarkable calculators these days. The last one I used was the venerable HP-41.

OpenStudy (zarkon):

I have a TI-Nspire CX CAS

OpenStudy (anonymous):

Interesting, thanks. I haven't owned a calculator in probably 15 years, since I'm usually never far from a computer, and it takes very little time to write a little C code to do anything more complex than 1 + 1. But the ability to easily do arithmetic with more than 15 digits precision is intriguing..

OpenStudy (zarkon):

whenever I need to program something I use MatLab